05-06 October 2008 On-sky TCS SW testing, AGw recalibration (N.Ageorges,W.Seifert, et al)

Observers: D. Thompson, LUCIFER folks
Instrument Support: Tim Shih
Software Support: M De La Pena in Tucson
Telescope Operator: A. Ceranski

Summary

Initial recalibration of the WFS worked. The system will close the loop and collimate the telescope. Found and fixed a couple of bugs, others being worked on (Issue Trak#1688-1690).

Details

  • Cleaned up /tmp directory on lbtmu107
    (moved images to subdirectories, nothing deleted).
  • Verified AG and WFS camera are alive, seeing scattered light in dome

Correct pointing for thermal drifts
IE went from -23.17 to -67.50, CA went from -28.96 to -26.70
05:30 UT IE=-82.18 CA=-25.66
06:45 UT IE=-77.60, CA=-15.60

First presets failed - OSS returning an error (No focal station defined) even though M3 is not being used. Paul is fixing it so that it does not do this. Acquisition images were obtained, but there was no star visible on the detector. Shortly after that we had to close for high humidity (UT2:20).

After Paul's fix...we found a bug in OSS in that the TRACK mode presets succeed but LUCIFER gets the same OSS message (this time not an error). It is now fixed (03:58).

Also found the SFP coordinates in the acquisition images at the beginning of the night are flipped in X from where they should be (needs to stay the same until the IAA/PAI issue is resolveD). The flip was removed in PCS, Michele is put it back and we confirmed it is correct by sending a preset inside the dome and verifying the "PROBSFPX" and "PROBSFPY" keywords in the header of the acquisition image.

Re-opening 05:16UT

Update pupil center on WFS

  • Preset to on-axis source, GUIDE mode, PA=0. Take WFS image
  • Evaluate center of pupil
  • Update file(s) as per D.Miller instructions

John is following Doug's suggestions... Center on first wfs image: 401,223

The IDL wavefront fitting code seems to be forcing the center of the projected pupil to be centered on one of the lenslets. However, the current AGW alignment (after replacing the camera) is clearly between two lenslets. GCS seems to be sending largely similar Zernikes.

Center in WFS program now at: 399,220 Sending preset in ACTIVE mode to confirm, failed because of HBS failure (twice!) WFS in GCS seems to be working pretty well!

Update AGw hotspot

  • Collimate "By-Eye"?
  • Preset to on-axis source, GUIDE mode, PA=0. Take WFS image
  • Preset to same source, GUIDE mode, PA=180. Take WFS image
  • Evaluate tip-tilt offset, update hotspot coords on guider to center on WFS
  • Repeat until good (<0.1 arcsec?)

First presets failed - OSS returning an error (No focal station defined) even though M3 is not being used. Paul is fixing it so that it does not do this. Acquisition images were obtained, but there was no star visible on the detector. Shortly after that we had to close for high humidity (UT2:20).

After Paul's fix...we found a bug in OSS in that the TRACK mode presets succeed but LUCIFER gets the same OSS message (this time not an error). It seems now to be fixed (03:58).

Also found the SFP coordinates in the acquisition images at the beginning of the night are flipped in X from where they should be (needs to stay the same until the IAA/PAI issue is resolveD). The flip was removed in PCS, Michele is put it back and we confirmed it is correct by sending a preset inside the dome and verifying the "PROBSFPX" and "PROBSFPY" keywords in the header of the acquisition image.

SX M1 Active Optics corrections at EL=71 deg:
X=-0.799 Y=+0.850 Z=+0.258 RX=9.17 RY=8.58

Check ACTIVE mode at lower ELEVATION

Preset to source at EL~45. Failed to find source on acq image...offset 22 arcsec! We need a new pointing model! Correct IE/CA. Preset succeeds. After allowing on-axis wfs to converge we preset in track mode to look at the full LUCIFER image. The stars were not very round.

SX M1 Active Optics corrections at EL=46 deg:
X=-0.320 Y=+0.692 Z=+0.183 RX=7.4 RY=3.4

07:45UT Closed again for rising humidity! Not yet done... ...and it won't be tonight. 4:45AM and the humidity is still 100%. Darn it!

*Maybe tomorrow:*
Update AGw on-axis position

  • Preset to on-axis source, ACTIVE mode, PA=0. Take WFS image
  • Preset to same source, GUIDE mode, PA=180. Take WFS image
  • Evaluate tip-tilt offset, update hotspot coords on guider to center on WFS
  • Repeat until good (<0.1 arcsec?)

07Oct08 DLM Moved GCS wfscimages to /Repository/AGW_Data/081006w_gcs and all GCS guiderimages to /Repository/AGW_Data/guider_081006-07

TCS verification of new functionality

Data for new M1 Collimation Model

-- DavidThompson - 05 Oct 2008
Topic revision: r9 - 08 Oct 2008, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback